OpenStudy (anonymous):

How do I find the Laplace transform of the following function (step functions used) f(t) = 3sin(pi*t)*u(t-2) - 3sin(pi*t)*u(t-3) ?

OpenStudy (turingtest):

check out example 2, I have to go... http://tutorial.math.lamar.edu/Classes/DE/StepFunctions.aspx

OpenStudy (anonymous):

not really helping...

OpenStudy (mr.math):

I think it's easier to use the definition here.

OpenStudy (anonymous):

how would i do that?

OpenStudy (anonymous):

anyone??????????????

OpenStudy (anonymous):

MY WAY \[3sin(\pi t)*u(t-2) - 3sin(\pi t)*u(t-3)\] \[3sin(\pi (t-2)+2\pi)*u(t-2) - 3sin(\pi (t-3)+3\pi)*u(t-3)\] \[3sin(\pi (t)+2\pi)u(t-2) - 3sin(\pi (t)+3\pi)u(t-3)\] \[3sin(\pi (t))u(t-2) + 3sin(\pi (t))u(t-3)\] \[3sin(\pi (t))u(t-2)= e^{-2s}(3\frac{\pi}{s^2+\pi^2})\] \[3sin(\pi (t))u(t-3)= e^{-3s}(3\frac{\pi}{s^2+\pi^2})\] \[e^{-2s}(3\frac{\pi}{s^2+\pi^2}+e^{-3s}(3\frac{\pi}{s^2+\pi^2})\]

OpenStudy (anonymous):

thanks!